OpenStudy (anonymous):

what are the asymptotes of f(x)=tan(2x-pi) from x=pi/2 to x=3pi/2?

OpenStudy (anonymous):

@Kainui

OpenStudy (ribhu):

put Cos(2x-pi) = 0

OpenStudy (ribhu):

ten calculate in the given range

OpenStudy (ribhu):

then...

OpenStudy (ribhu):

@Patsfan12 atleast you can ask your doubt or anything.

OpenStudy (anonymous):

Pi/2

OpenStudy (ribhu):

no

OpenStudy (ribhu):

\[2x-\pi = (2n+1)\pi/2, x= 3\pi/4, 5\pi/4\] in the given range.

OpenStudy (anonymous):

My answer choices are: X=3pi/4,x=5pi/4 X=0x=pi, x=2pi X=0,x=pi/4 X=pi/2, x=3pi/2

OpenStudy (anonymous):

So the first choice!

OpenStudy (ribhu):

yeah. and i am a decent guy you took me wrong

OpenStudy (anonymous):

Right you are again

OpenStudy (ribhu):

but u've blocked me.

OpenStudy (anonymous):

I unblocked you

OpenStudy (ribhu):

really tell me would u talk?

OpenStudy (ribhu):

lets be friends

OpenStudy (ribhu):

i will try to help you the way i can.